OpenStudy (anonymous):

If the linear equation 3x = 9 − 2y is written in standard form, what is the value of B?

OpenStudy (ghazi):

where is B??

OpenStudy (anonymous):

i dont know

OpenStudy (anonymous):

you have to put the equation in standard form but i don't know how to

OpenStudy (chihiroasleaf):

3x+2y=9 is the standard form

OpenStudy (anonymous):

Standard form is Ax+By=C without fractions and where A is non-negative

OpenStudy (anonymous):

okay...

OpenStudy (ghazi):

\[3x+2y=9......\frac{ 3x }{ 9 }+\frac{ 2y }{ 9 }=1\] now yur B= 9/2

OpenStudy (anonymous):

so the answer is 9/2????

OpenStudy (anonymous):

@chihiroasleaf nailed it... What is B in his statement?

OpenStudy (chihiroasleaf):

@ghazi should we make the constant become 1 to be the standard form?

OpenStudy (anonymous):

no fractions allowed

OpenStudy (chihiroasleaf):

@EulerGroupie I don't know what is the B, it's not clear on the question if B is the constant of 'y' then B = 2

OpenStudy (ghazi):

yes....your standard equation is like\[\frac{ x }{ A }+\frac{ y }{ B }=1\]

OpenStudy (anonymous):

Sorry, but no. It is Ax+By=C

OpenStudy (anonymous):

so what would the answer be then?

OpenStudy (anonymous):

2

OpenStudy (ghazi):

sorry it's 9/2

OpenStudy (chihiroasleaf):

I think @EulerGroupie is correct you can read this http://www.algebralab.org/studyaids/studyaid.aspx?file=algebra1_5-5.xml so the answer is 2

OpenStudy (anonymous):

I deal with this all the time... don't trust me... try a web search.

OpenStudy (anonymous):

are you being serious or sarcastic?^

OpenStudy (anonymous):

Standard equation for a plane Ax+By+Cz=D

OpenStudy (anonymous):

No, serious. You don't have to trust me. I did a web search, it came up immediately.

OpenStudy (anonymous):

so the answer is 2?

OpenStudy (chihiroasleaf):

yes... :)

OpenStudy (ghazi):

yes ...it was my fallacy @EulerGroupie is right

OpenStudy (anonymous):

No problems @ghazi. I know that you have amazing skills. I tutor at a community college and there are some questions that I field about 10 times per semster and this is one of them. I personally think the form is almost useless. It works well to find x and y intercepts, otherwise I like slope intercept or point slope better.

OpenStudy (ghazi):

exactly ....actually what i wrote is for finding slope....

OpenStudy (anonymous):

okay thanks guys! i appreciate it (:

OpenStudy (anonymous):

You're welcome.

OpenStudy (ghazi):

not slope it's intercept

OpenStudy (chihiroasleaf):

you're welcome :)

OpenStudy (anonymous):

Cool, an analytic form (like conic sections) of a line that states the intercepts. I like it @ghazi. That is one of the reasons that I like this site... new perspectives.